# Rapid Bayesian identification of sparse nonlinear dynamics from scarce and noisy data

### v1

@@ 4.1 Scaling of likelihood and Occam factor with $N$ and $\boldsymbol{\beta}$
Conversely, figure 11$(a,c,e)$ demonstrates that the Occam factor assumes greater significance in model selection when noise is large and data is scarce. The Occam factor serves two crucial functions in model selection. First, it penalises models with too many terms, as it scales with $M$. Second, it quantifies the variance of the estimated parameter, and discards terms that have small variances in their coefficients, as this indicates that the corresponding model is not robust (or too sensitive to small changes in the parameter values). In performing the above two functions, the rigorously derived Occam factor excludes terms that incidentally fit the data, ensuring the robustness of the estimated model.
4.2 Limitation and use cases of Bayesian-SINDy
Although Bayesian-SINDy has shown greater resilience than the original STLS-based SINDy in accurately identifying the proper dynamical equations from noisy and scarce datasets, there is still much room for improvement.
For example, one might find the noise level in the nonlinear (cubic) oscillator example in §3.2 low compared to other examples. In fact, none of the three algorithms were able to recover the correct equations upon further increases in noise level. This is because, in the presence of highly nonlinear cubic terms, the noise distribution can become highly skewed and deviate from a Gaussian distribution upon an increasing noise-to-signal ratio. As we have stressed in (9) and in §2.2, the Gaussian distribution is only strictly valid if the noise is truly Gaussian and the model consists of linear terms only. However, it serves as a good approximation in most cases, as shown empirically by the other examples. Alternatively, when the noise distributions are no longer tractable or well approximated by analytical methods, one can resort to sampling-based techniques to better capture the accurate noise distribution. For example, recent work by Hirsh et al. [31] has demonstrated the possibility of further improving noise robustness and data efficiency by accurately capturing distributions through sampling techniques such as MCMC. However, this improvement comes at the expense of costly sampling. For example, it took days of computation to recover the Lotka-Volterra equation from the Hudson Bay Lynx-Hare population dataset using the MCMC-based method in [31], compared to the sub-second computation by Bayesian-SINDy.
@@ 4.2 Limitation and use cases of Bayesian-SINDy
Consequently, the choice of which approach to adopt depends on the specific application and circumstances. If computing time and cost are not significant constraints, sampling-based methods such as those suggested by [30] or [31] should give more accurate models than Bayesian-SINDy. Nonetheless, Bayesian-SINDy is a valuable tool for conducting rapid preliminary investigations. In situations that demand rapid learning, such as system identification in control, using a sampling-based method would not be practical. Instead, Bayesian-SINDy could be a cost-effective approach to extracting symbolic models from noisy and limited data in a fast and reliable manner.
@@ 4.3 Application of Bayesian-SINDy in active learning
which is based on the evidence $p(\mathbf{t}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})$. In a sampling-based technique, the evaluation of $p(\mathbf{t}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})$ already requires the integration of the posterior in the parameter space. The evaluation of $H$, then, would have required further integration in the $\mathbf{t}$-space, which would be prohibitively expensive if a sampling technique were used. Fortunately, the expensive evaluation can be avoided if we accept the Gaussian approximation used throughout this work. Then, the analytical tractability of the Gaussian distribution allows for $H$ to be rewritten explicitly as
@@
after substituting (14-15) into (26). Here, $\mathbf{C}$ is the covariance matrix of the evidence.
@@
In real-life applications such as system identification for control, one can imagine using $H$ as part of an objective function for optimal control. By balancing the objective of learning more about the dynamics governing the system (i.e. exploration) with the objective of achieving certain target state(s) (i.e. exploitation), one can develop the current framework into a fully automated active learner that can learn about the system and drive it to a certain state in a very rapid manner.
@@ 5 Conclusion and future work
In this work, we have presented a unifying Bayesian framework to compare Bayesian-SINDy with the original STLS-based SINDy. By grounding STLS in the Bayesian framework, we revealed how STLS can be a robust model selection method in the high-data low-noise limit due to the dominant contribution of the likelihood to the evidence. However, in the low-data high-noise limit, the Occam factor becomes more important, and Bayesian-SINDy’s ability to penalise overfitting becomes crucial. The Bayesian framework also provides a more rigorous justification for selecting more parsimonious models than SINDy. Instead of selecting models by heuristically tweaking hyperparameters towards the Pareto front, the Bayesian framework formally defines the Occam factor, resulting in a more principled approach to model selection. This results in the Bayesian-SINDy algorithm, which provides a more concrete theoretical foundation for current SINDy-based approaches and a more robust framework for further extensions. Along with other recent publications highlighting the Bayesian perspective on SINDy, such as [33]’s demonstration of the convergence of Ensemble-SINDy to Bayesian statistics, we believe that many current SINDy extensions can benefit from the Bayesian framework to improve robustness in noisy and scarce datasets.
To demonstrate how rigorously defined Occam factor helps model selection in the low data regime, we substituted STLS with sparse Bayesian learning to create the Bayesian-SINDy algorithm. Notably, in all examples considered, the Bayesian-SINDy algorithm was able to recover the intended equations most of the time in hundreds of data points or fewer, demonstrating its data efficiency. We have also compared Bayesian-SINDy to SparseBayes, which is the foundation of previous Bayesian extensions to SINDy, and demonstrated the importance of considering how noise propagates through the nonlinear candidate functions and the variation of noise level over time.
@@
Aside from improving the approximation of the noise distribution, much work can be done to expand the current methodology. For example, while this work has primarily focused on retrieving desired equations from homogeneous datasets where data are synthesized from the same unknown parameters, our methodology can also combine multiple datasets that share the same underlying dynamics but with different unknown parameters. The original SINDy-based method can train parameterised models from datasets of different parameters if the parameters are known or measured as part of the dataset. However, fixing or knowing all parameters in real-world data is typically difficult or impossible to achieve. For example, in modelling cell motility, a population of cells may share a common model for their motility, yet each cell might be parameterised differently and it is difficult to accurately measure each parameter before learning the model. Because the evidence is derived by marginalising the parameters, it is independent of the specific parameter value and provides a mathematical framework to combine multiple datasets with varying unknown parameters.
@@
Lastly, it would also be interesting to benchmark Bayesian-SINDy with other recent advancement in SINDy, such as the ensembling technique [12] and the denoising method based on automatic differentiation [38]. In particular, recent work has established open standard for benchmarking different sparse system identification methodologies [39]. While we do not expect the Bayesian-SINDy to be the most noise-robust or most accurate in all scenarios, its relatively cheap computational cost and excellent data efficiency make it more suitable for real-time applications such as model identification in control.
Appendix A. Selection of hyperparameters values
@@
Cubic Oscillator $1$ $\phi=(t^{2}-1)^{4}$, 6 points

### v2

@@ 4.1 Scaling of likelihood and Occam factor with $N$ and $\boldsymbol{\beta}$
Conversely, figure 11$(a,c,e)$ demonstrates that the Occam factor assumes greater significance in model selection when noise is large and data is scarce. The Occam factor serves two crucial functions in model selection. First, it penalises models with too many terms because it scales with $M$. Second, it quantifies the variance of the estimated parameters, and discards terms that have small variances in their coefficients because this indicates that the corresponding model is not robust (or is too sensitive to small changes in the parameter values). In performing the above two functions, the rigorously derived Occam factor excludes terms that incidentally fit the data, ensuring the robustness of the estimated model.
4.2 Limitations and applications of Bayesian-SINDy
One of the key contributions of this study is the development of a computationally efficient estimation of the noise variance, which is critical in calculating the evidence Bayesian-SINDy optimizes for during the model selection process. Mutliple approximations were made as a trade-off between computational efficiency and accuracy, which may limit the applicability of Bayesian-SINDy despite its greater resilience than the original STLS-based SINDy in accurately identifying the proper dynamical equations from noisy and scarce datasets. Therefore, here we provide a review of the approximations made, their limitations, and potential remedies.
Gaussian noise
The noise model for $\boldsymbol{\epsilon}_{x}$ is assumed to be Gaussian, which is a common assumption in many statistical methods and is chosen in this study mainly for its tractability. Real-world data may not, however, follow a Gaussian distribution. Nonetheless, the framework for the noise estimation laid out in §2.1 is built upon a Taylor expansion of moments, which is not restrictive to Gaussian noise. Therefore, this framework can be extended to other noise distributions by modifying the noise estimation procedure in §2.1 to account for the different moments of the noise distribution.
Noise correlation in time
One major assumption in the noise model is that the noise is uncorrelated in time. This assumption was made mostly to speed up computation, but is not neccessary. In fact, one can easily relax the assumption by replacing the inverse variance $\boldsymbol{\beta}$ with the full inverse covariance matrix $\mathbf{B}$. The downside of this approach is that the computational cost of the regression procedure would increase from $\mathcal{O}(N)$ to $\mathcal{O}(N^{3})$, which may be prohibitive for large datasets.
Noise correlation between terms
Another assumption used to speed up computation was that the noise is uncorrelated between terms. If the noise correlation between terms is strong, the noise estimation may not be accurate. Unfortunately, there is no remedy to this issue without incurring significant computational cost, as there are combinatorially many correlations between each term to be considered while the algorithm selects the best model. However, one can always check the validity of the noise model by comparing the estimated noise variance with MCMC a posteriori.
Estimating the noise distribution from the nonlinear functions
As mentioned in §2.1, as the noise propagates through the nonlinear functions in $\boldsymbol{\Theta}(\mathbf{X})$, the resulting noise distribution is non-Gaussian. While we use a Taylor expansion to approximate the resulting variance, this estimation is only valid when the noise is relatively small such that higher order terms can be neglected.
Figure 12: Distribution of the total noise $\boldsymbol{\epsilon}$ for $x_{2}$ at $t=2.5$ in the cubic oscillator example with $\Delta t=0.1$. Here, the distribution is created from 10000 samples with noise level $(a)$ $\sigma_{x}=0.05$ and $(b)$ $\sigma_{x}=0.5$. The histogram shows the distribution from data, while the red line shows the Gaussian estimation using variance $\boldsymbol{\sigma}^{2}$ from (8). Notice that the distribution is skewed and deviate from the Gaussian distribution when the noise is too big.
For example, in the nonlinear oscillator case, if one further increases the noise, none of the three algorithms will be able to recover the correct equations. This is because, in the presence of highly nonlinear cubic terms, the noise distribution can become highly skewed, as shown in 12, breaking the variance and mean approximation from the Taylor expansion of moments. To capture the skewed distribution, one must extend the algorithm to consider higher order moments, incurring significant computational cost. Given the restriction in noise level under this approximation, a user may want to gauge the validity of this approximation. One way to probe its validity a priori is to sample around the data with the given noise level and pass the sample through the nonlinear functions. By comparing the skewness and conformity of the resulting distribution with what would have been expected from the approximation (i.e., a Gaussian distribution), one can determine if the proposed algorithm provides suitable approximation of the noise distribution. If the application is not time-critical, it is also a good practice to check the generated model against MCMC sampling a posteriori to ensure the noise model’s validity.
Gaussian approximation of the likelihood
In §2.1, after propagating the noise through the transformation in (6) by tracing the transformation on the noise variance $\boldsymbol{\sigma}^{2}$, we arrive at the total noise $\boldsymbol{\epsilon}$ in (11) with the knowledge of its variance. We then approximate the likelihood to be Gaussian with variance $\boldsymbol{\sigma}^{2}$. This final Guassian approximation is what enables the proposed regression-based algorithm and SparseBayes to be fast. This approximation can be justified by the fact that, in the context of performing type-II maximum likelihood optimisation, a Laplace’s approximation to the evidence is usually good enough to rank models in the correct order. The Laplace’s approximation only requires the knowledge of the Maximum a Posteriori (MAP) probability and the Hessian around the MAP point, which can be approximated as the variance of the posterior, and is fairly accurate as long as the posterior is unimodal and does not have a heavy tail.
Applications suitable for Bayesian-SINDy
@@ 4.2 Limitations and applications of Bayesian-SINDy
Although the computational efficiency of Bayesian SINDy requires many approximations, the algorithm still selects the correct models robustly and accurately in many cases, as demonstrated empirically in §3. However, when the noise distributions are no longer tractable or well approximated by the proposed method, e.g. when the signal-to-noise ratio is too low or the noise is too skewed, it is recommended to resort to sampling-based techniques to better capture the accurate noise distribution. Recent work by Hirsh et al. [32] has demonstrated the possibility of further improving noise robustness and data efficiency by accurately capturing distributions through sampling techniques such as MCMC. However, this improvement comes at the expense of costly sampling. For example, it took hours of computation to recover the Lotka-Volterra equation from the Hudson Bay Lynx-Hare population dataset using the MCMC-based method in [32], compared to the sub-second computation by Bayesian-SINDy. Even if a further a posteriori check is performed using MCMC, the overall computation time would still be significantly lower than Hirsh’s method [32]. Consequently, the choice of which approach to adopt depends on the specific applications and circumstances. If computing time and cost are not significant constraints, sampling-based methods such as those suggested by [31] or [32] should give more accurate models than Bayesian-SINDy. Conversely, if time or computational cost is a major concern, Bayesian-SINDy is a valuable tool for conducting rapid preliminary investigations. In situations that require rapid learning, such as system identification in control, using a sampling-based method would not be practical. Instead, Bayesian-SINDy could be a cost-effective approach to extracting symbolic models from noisy and limited data in a fast and reliable manner.
@@ 4.3 Application of Bayesian-SINDy in active learning
which is based on the evidence $p(\mathbf{y}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})$. In a sampling-based technique, the evaluation of $p(\mathbf{y}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})$ already requires the integration of the posterior in the parameter space. The evaluation of $H$, then, would have required further integration in the $\mathbf{y}$, which would be prohibitively expensive if a sampling technique were used. Fortunately, the expensive evaluation can be avoided if we accept the Gaussian approximation used throughout this study. Then, the analytical tractability of the Gaussian distribution allows for $H$ to be rewritten explicitly as
@@
after substituting (13-14) into (25). Here, $\mathbf{C}$ is the covariance matrix of the evidence.
@@
In real-life applications such as system identification for control, one can imagine using $H$ as part of an objective function for optimal control. By balancing the objective of learning more about the dynamics governing the system (i.e. exploration) with the objective of achieving certain target state(s) (i.e. exploitation), one can develop the current framework into a fully automated active learner that can learn about the system and rapidly drive it to a certain state.
@@ 5 Conclusion and future work
In this study, we have presented a unifying Bayesian framework to compare Bayesian-SINDy with the original STLS-based SINDy. By grounding STLS in the Bayesian framework, we have revealed how STLS can be a robust model selection method in the high-data low-noise limit due to the dominant contribution of the likelihood to the evidence. However, in the low-data high-noise limit, the Occam factor becomes more important, and Bayesian-SINDy’s ability to penalise overfitting becomes crucial. The Bayesian framework also provides a more rigorous justification for selecting more parsimonious models than the STLS-based SINDy. Instead of selecting models by heuristically selecting hyperparameters at the Pareto front, the Bayesian framework formally defines the Occam factor, resulting in a more principled approach to model selection. This results in the Bayesian-SINDy algorithm, which provides a more concrete theoretical foundation for current SINDy-based approaches and a more robust framework for further extensions. Along with other recent publications highlighting the Bayesian perspective on SINDy, such as the demonstration of the convergence of Ensemble-SINDy to Bayesian statistics found in [34], we conclude that many current SINDy extensions could benefit from the Bayesian framework to improve robustness in noisy and scarce datasets.
To demonstrate how a rigorously defined Occam factor helps model selection in the low data regime, we substituted STLS with sparse Bayesian learning to create the Bayesian-SINDy algorithm. Notably, in all examples considered, the Bayesian-SINDy algorithm was able to recover the intended equations most of the time in hundreds of data points or fewer, demonstrating its data efficiency. We have also compared Bayesian-SINDy to SparseBayes, which is the foundation of previous Bayesian extensions to SINDy, and demonstrated the importance of considering how noise propagates through the nonlinear candidate functions and the variation of noise level over time.
@@
Aside from improving the approximation of the noise distribution, much work can be done to expand the current methodology. For example, while SINDy can be extended to find PDEs [10], Bayesian-SINDy can also be extended in the same way. Another possible extension is the possibility to combine multiple dataset with different parameters. While this study has primarily focused on retrieving desired equations from homogeneous datasets in which data are synthesized from the same unknown parameters, our methodology can also combine multiple datasets that share the same underlying dynamics but with different unknown parameters. The original SINDy-based method can train parameterised models from datasets of different parameters if the parameters are known or measured as part of the dataset. However, fixing or knowing all parameters in real-world data is typically difficult or impossible to achieve. For example, in modelling cell motility, a population of cells may share a common model for their motility, yet each cell might be parameterised differently and it is difficult to accurately measure each parameter before learning the model. Because the evidence is derived by marginalising the parameters, it is independent of the specific parameter value and provides a mathematical framework to combine multiple datasets with varying unknown parameters.
@@
Lastly, it would also be interesting to benchmark Bayesian-SINDy with other recent advancement in SINDy, such as the ensembling technique [12] and the denoising method based on automatic differentiation [42]. In particular, recent work has established an open standard for benchmarking different sparse system identification methodologies using chaotic systems [43]. While we do not expect the Bayesian-SINDy to be the most noise-robust or most accurate in all scenarios, its relatively cheap computational cost and excellent data efficiency make it more suitable for real-time applications such as model identification in control. Moreover, given Bayesian-SINDy’s rigorous justification for model selection under the Bayesian framework, it would be interesting to investigate how this theoretical justification may or may not help its performance in the benchmark compared to other methodologies.
Acknowledge, contributino and data
The Bayesian-SINDy method is available at https://github.com/llfung/B-SINDy. Contributions: L.F.: conceptualisation, formal analysis, methodology, investigation, writing, visualization; U.F.: supervision, validation, review and editing; M.P.J.: conceptualisation, investigation, supervision, review and editing. L.F. acknowledge the support of Research Fellowship from Peterhouse, Cambridge.
Appendix A. Formula for the library of variance
To calculate the variances of each column of the library of polynomial terms, we follow the method of Taylor expansion for the moments of functions of random variables [35].
Given a random variable $X$ with Gaussian distribution $X\sim\mathcal{N}(\mu,\sigma_{x}^{2})$, the moments of the function $f(X)$ can be calculated by the Taylor expansion of the function around the mean $\mu$. In particular, the power function $f(X)=X^{n}$ has the following moments:
$E[X^{n}]=\sum_{i=0}^{n}\frac{\sigma_{x}^{i}}{i!}\frac{n!}{(n-i)!}\mu^{n-i}a_{i},$ (27)
where the coefficients $a_{i}$ are given by
$a_{i}=\begin{cases}\prod_{j=1}^{i/2}(2j-1)&\text{if }i\text{ is even}\\ 0&\text{if }i\text{ is odd}\end{cases}$ (28)
and
$\sigma^{2}_{x^{n}}=\text{Var}[X^{n}]=E[X^{2n}]-(E[X^{n}])^{2}.$ (29)
To calculate the variance of the weighted sum or product of two independent random variables $X_{1}$ and $X_{2}$, we use the formula
$\sigma^{2}_{ax_{1}+bx_{2}}=\text{Var}[aX_{1}+bX_{2}]=a^{2}\text{Var}[X_{1}]+b^{2}\text{Var}[X_{2}]$ (30)
and
$\sigma^{2}_{x_{1}x_{2}}=\text{Var}[X_{1}X_{2}]=E[X_{1}^{2}]E[X_{2}^{2}]-(E[X_{1}X_{2}])^{2}.$ (31)
Appendix B. Selection of hyperparameters values
@@
Cubic Oscillator $1^{2}$ $\phi=(t^{2}-1)^{4}$, 6 points